OpenStudy (anonymous):

Point K is graphed on a number line at 3. Point L is 11 units away from point K on the number line. What could be the coordinates of point L? Choose all answers that are correct. A. –14 B. –8 C. 7 D. 14

OpenStudy (anonymous):

3 + 11 = 3 - 11 = both answers are 11 units from 3

OpenStudy (anonymous):

-8 and 14

OpenStudy (anonymous):

Oops.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

thank chu

OpenStudy (anonymous):

wait so its d and b

OpenStudy (anonymous):

both

OpenStudy (anonymous):

ok